OpenStudy (anonymous):

can anyone help me with this problem please, 2^x^2=(8)2x*1/256

OpenStudy (anonymous):

please clearly write the equation i can't understand it.

OpenStudy (anonymous):

it is 2 raised to x^2= 8^2x multiplied by 1/256 thanks for your time.

OpenStudy (anonymous):

\[2^{2x}=8^{2x}/256\] since \[256=2^{8}\] therefore \[2^{2x}=8^{2x}/2^{8}\] since \[8^{2x}=(2^{3})^{2x}=2^{6x}\] hence \[2^{2x}=2^{6x-8}\] therefore bases are same & powers are only solved... 2x=6x-8 4x=8 x=2